OpenStudy (anonymous):

Classify the conic section and write its equation in standard form. 1) y2 - 8y - 8x = -24

OpenStudy (anonymous):

going to be a parabola what is your version of "standard form"?

OpenStudy (anonymous):

maybe it is \((y-k)^2=4p(x-h)\)

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

it says "write in standard form" so to do this we must know exactly what "standard form" means, otherwise the question makes no sense

OpenStudy (anonymous):

i think that is the standard form

OpenStudy (anonymous):

ok then we can do it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[y^2 - 8y - 8x = -24\] add \(8x\) get \[y^2-8y=4x-24\] then "complete the square to get \[(y-4)^2=4x-28+16\] or \[(y-4)^2=4x-12\] factor on the right and get \[(y-4)^2=4(x-3)\]

OpenStudy (anonymous):

thanks can u help with 2 more just like this one ??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x2 + y2- 4x + 2y - 11 = 0

OpenStudy (anonymous):

for this one you have to complete the square as well since the square terms are \(x^2\) and \(y^2\) it is a circle

OpenStudy (anonymous):

standard form will be \[(x-h)^2+(y-k)^2=r^2\] start by grouping terms \[x^2 + y^2- 4x + 2y - 11 = 0\] \[x^2-4x+y^2+2y=11\]

OpenStudy (anonymous):

then complete the square for both \[(x-2)^2+(y+1)^2=11+4^2+1^2\] \[(x-2)^2+(y+1)^2=29\]

OpenStudy (anonymous):

center of the circle is \((2,-1)\) and radius is \(\sqrt{29}\) what you need to know in order to do these is how to complete the square

OpenStudy (anonymous):

kay thanks .

OpenStudy (anonymous):

4x2 + y2 - 8x - 8 = 0

OpenStudy (anonymous):

this one is an ellipse because the coefficient of the \(x^2\) and \(y^2\) terms are different

OpenStudy (anonymous):

standard form is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]

OpenStudy (anonymous):

\[4x^2 + y^2 - 8x - 8 = 0\] \[4x^2-8x+y^2=8\]\[4(x^2-2x)+y^2=8\] \[4(x-1)^2+y^2=8+4=12\]then divide by 12 and get \[\frac{(x-1)^2}{3}+\frac{y^2}{12}=1\]

OpenStudy (anonymous):

okay i really appreciate i have one more 4x2 - 25y2 - 100y = 0

OpenStudy (anonymous):

this one a hyperbola because of the \(-25y^2\)

OpenStudy (anonymous):

did you copy it correctly ?

OpenStudy (anonymous):

4x2 - 25y2 - 100y = 0 yes

OpenStudy (anonymous):

ok then it is nothing

OpenStudy (anonymous):

so its not a standard form?

OpenStudy (anonymous):

ok wait i made a mistake, hold on

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[4x^2 - 25y^2 - 100y = 0 \]\[4x^2-25(y^2+4)=0\] \[4x^2-25(y+2)^2=-100\] \[25(y+2)^2-x^2=100\] \[\frac{(x+2)^2}{4}-\frac{x}{100}=1\]

OpenStudy (anonymous):

typo on the last line, should be \[\frac{(y+2)^2}{4}-\frac{x}{100}=1\]

OpenStudy (anonymous):

thank you !